User Tools

Site Tools


This wiki is not maintained! Do not use this when setting up AuScope experiments!

Meeting 23rd June

Chair: Jonny

Present: Mas, Arwin, Jim, Lucia, Jamie, Liz, Patrick, Jesse

Apologies: Tiege, Lucas


General Issues:

* Fill in the meeting chair roster
* New queen couch/not-bed in the lounge/not-bedroom.
* Baby monitor now available
  • Hobart 12m are out for a while (~month+)
  • Triple check when your shifts are!

Handover Issues:

* All 3 12m seem to have many erc and fs lockups lately
  - Can be caused by dodgy modules (happened a bit lately) 

* Hb: unable to change modules - R4795

* Ke: clock drifts/DBBC issues
  - Fixed by power cycling/reconfiguring DBBC. Make a note/report if you have to do this.
  - Can be caused by power fluctuations. 
* Ke: weather command 'wx' locking up (returning identical values?)
  - Jamie fixing/fixed? it
  - Issues with javascript that pulls the data from the geoscience aus website?
  - anyone interested in solving this, talk to Jamie and get a choccie if you fix it.
  - Make sure to note in the end of exp. notes that the wx values may not be correct
* Ke: network down a lot
  - Patrick noticed Telstra were having adsl issues in the area
* Ke: Interior light not working
  - Makes reading MK5 labels difficult
  - May be good to ghetto up some sort of small light 

* Yg: Drives getting stuck a lot
   - Elevation drives getting stuck,
   - a fix that has worked: turn drives off, wait, wait more, press operate
  • Jim doing some training for Mas and Patrick for the 26m, probably
    1. Tuesday out at Mt. P. 10amish.
/home/www/auscope/opswiki/data/pages/operations/agenda2017_06_23.txt · Last modified: 2017/06/23 05:25 by Jonathan Rogers